# Seeding with Differentially Private Network Information

### v2

@@ 3 Differentially Private Seeding Algorithms
In this section, we provide two DP seeding algorithms that satisfy the DP requirements either centrally or locally. First in Subsection 3.1, we use an exponential mechanism that ensures the output of the seed selection is centrally DP. Next in Subsection 3.2, we use the randomized response algorithm to achieve local DP according to the criteria set forth in Definition 2.4. LABEL:thm:exp:acc and LABEL:thm:rr:acc provide our sample complexity bounds in each case to achieve $\epsilon$-ISDP with a $(1-1/e)\mbox{\sf OPT}-\alpha n$ performance guarantee. In particular, we observe that the sample complexity bound for achieving local DP is much harsher at $O(k^{3}\epsilon^{-2k^{2}}\ln n/\alpha^{2})$ compared to $O(k\ln n/(\alpha\epsilon))$ for central DP. This exponential separation of sample complexity dependence on $k$ is reflected in our simulation results in Section 5, where exponential mechanism shows a faster improvement in performance with increasing number of influence samples and its output has a significantly higher expected spread size when the privacy budget ($\epsilon$) is low.
@@ 3.1 Central DP with Exponential Mechanism
In this section, we design a DP seeding algorithm by using exponential mechanism and modifying Algorithm 1. To protect privacy, instead of Algorithm 2, we use an exponential mechanism that differentially outputs a soft max (LABEL:thm:exp:utility). Formally, given $\epsilon>0$ and integer $k$ we define a seeding algorithm $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ that runs Algorithm 1 but replaces the single seed selection steps of Algorithm 2 by its exponential mechanism variant, Algorithm 3 $\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}},S)$. Note that Algorithm 3 is an exponential mechanism using $I_{\mathbf{x}}(\cdot|S)$ as the scoring function.
@@
Note that a naive algorithm that runs the exponential mechanism directly on all $k$ seeds at once is not efficient because the size of the output space is combinatorially large $\Omega(n^{k})$. To ensure our algorithm is efficient, we sequentially run the exponential mechanism $k$ times to estimate and choose seeds similar to the greedy algorithm of Eckles et al. (2022), reducing their time complexity to $O(nmk)$. LABEL:thm:exp:inf shows that our exponential mechanism for single seed selection, $\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}},S)$, is $\epsilon$-ISDP (Definition 2.4), and provides a tail bound on the approximation error of maximizing $I_{\mathbf{x}}$. Its proof in Section 7.3 follows from the facts that $I_{\mathbf{x}}$ is a submodular function, and the exponential mechanism in Algorithm 3 can be seen as a noisy maximum function which implements an approximate greedy step by LABEL:thm:exp:utility.
@@
Given graph ${\mathcal{G}}$ on $n\geq 2$ nodes and $k\leq n$, for any $0<\alpha\leq 1$ and $\epsilon>0$, $\mathcal{M}^{k,\epsilon}_{exp}(\cdot)$ that runs Algorithm 1 but replaces the single seed selection steps of Algorithm 2 by $\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}},S)$, its exponential mechanism variant in Algorithm 3, is $\epsilon$-ISDP and its output infects at least $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes with high probability, using at least $m=\max\{\frac{12}{\alpha\epsilon},\frac{9}{\alpha^{2}}\}k\ln n$ influence samples in $O(knm)$ run time.
Note that if we are not concerned with privacy and set $\epsilon\gg\alpha$, then the number of samples to achieve $\alpha n$ additive error is of order $\alpha^{-2}k\ln n$, which improves upon the $k^{2}(\ln n/\alpha)/\alpha^{3}$ bound of Eckles et al. (2022). Moreover, our result does not violate Theorem 3 of Eckles et al. (2022) that shows $\Omega(k^{2})$ influence samples are required to achieve $(1-1/e)\mbox{\sf OPT}-\alpha n$ performance bound for all $\alpha<(1-1/e)/k$, because our algorithm also needs $\Omega(k^{2})$ samples when $\alpha<1/k$.
@@ 3.2 Local DP with Randomized Response
We can achieve a stronger notion of privacy, local DP, by adding noise to the input data before running our (or any other) algorithms. Line 3 of Algorithm 4 adds the requisite amount of noise to achieve local $\epsilon$-ISDP in the input dataset ${\mathbf{x}}$ — flipping each entry of ${\mathbf{x}}$ independently with probability $\rho_{\epsilon}=1/(1+e^{\epsilon})$ to generate the noisy input $\tilde{{\mathbf{x}}}$. However, the added input noise will bias our estimate of each candidate seed’s quality; in particular, $\operatorname*{\mathbb{E}}_{\tilde{{\mathbf{x}}}}[I_{\tilde{{\mathbf{x}}}}(S|T)]$ is not necessarily equal to $I_{\mathcal{G}}(S|T)$ for a pair of subsets $S,T\subset V$ and we no longer have the benefit of applying Proposition 2.1 directly. Rather in Algorithm 5 we propose a post-processing of the noisy input data $\tilde{{\mathbf{x}}}$ to derive an unbiased estimate of a seed set’s quality. Importance of proper post-processing steps in the analysis of DP data are highlighted in other applications, e.g., census DP data release, where ignoring them can lead to biased decision-making and sub-optimal outcomes (Abowd et al. 2019, Ashmead et al. 2019, Zhu et al. 2022, 2021). The added variability of the estimators due to the DP noise can also degrade our estimation performance (even if not biased), which is a second concern in local DP algorithm design and opens a venue to explore alternative post-processing methods with potentially superior bias-variance tradeoff. In the remainder of this section, we first introduce Algorithms 4 and 5 and then offer LABEL:thm:rr:acc for their privacy and utility guarantees, addressing both concerns about the bias and variance of decision making with local DP influence samples.
@@
Given a set $S$, an influence sample $x\in\{0,1\}^{n}$, and its perturbed version $\tilde{x}$, consider two sets of nodes $S\cap x$ and $S\cap\tilde{x}$. By Proposition 2.1, the influence function $I_{\mathcal{G}}(S)$ can be estimated from the reachable set, $n\Pr[S\cap x\neq\emptyset]$. But to achieve local DP, we can only observe the perturbed, reachable set $S\cap\tilde{x}$. In other words, let $\tilde{z}_{S}:=|S\cap\tilde{x}|$ and ${z}_{S}:=|S\cap{x}|$ be two random variables. If we can estimate $\Pr[z_{S}=0]$ from our observed samples of $\tilde{z}_{S}$, then we can estimate the size of the reachable set from $I_{\mathcal{G}}(S)=n\Pr[z_{S}>0]$ — using Proposition 2.1 — and then run the standard greedy algorithm for seed selection based on the estimated values.
Now, we discuss how to estimate $\Pr[z_{S}=0]$ from samples of $\tilde{z}_{S}$. The main observation is that the probability mass function of $\tilde{z}_{S}$ is a linear function of the probability mass function of $z_{S}$. Note that $S\cap x$ is the set of elements in $S$ that is also in the influence sample $x$. Hence, $S\cap\tilde{x}$ can be partitioned into (1) the subset $(S\cap\tilde{x})\cap x$ of elements of $x$ that remain in $S$ after perturbation, and (2) the subset $(S\cap\tilde{x})\setminus x$ of the elements of $\tilde{x}$ that join $S$ after the perturbation. The size of the first subset is a binomial random variable $Bin(|S\cap x|,1-\rho)$, and the size of the second subset is $Bin(|S\setminus x|,\rho)$. Therefore, given $|S|=l$ for all $a$ and $b$, the probability of $\tilde{z}_{S}=b$ conditional on $z_{S}=a$ is $\Pr[\tilde{z}_{S}=a|z_{S}=b]=\Pr[Bin(b,1-\rho)+Bin(l-b,\rho)=a],$ for all $a,b=0,\dots,l$ where $Bin(b,1-\rho)$ and $Bin(l-b,\rho)$ are independent. Additionally, if we set matrix $C=C_{\rho,l}\in{\mathbb{R}}^{(l+1)\times(l+1)}$ be a likelihood function where $C(a,b)=C_{\rho,l}(a,b)=\Pr[\tilde{z}_{S}=a|z_{S}=b]$, then we have:
@@
We can write out each entry explicitly as follows:
@@
To see why, condition on $\ell$ of the $z_{S}=b$ nodes in $S\cap x$ being flipped, leaving $b-\ell$ nodes in $S\cap x\cap\tilde{x}$. To get $\tilde{z}_{S}=a$ nodes in $S\cap\tilde{x}$, the remaining $a-b+\ell$ nodes should come from perturbing the status of the $l-b$ nodes belonging to $S\setminus x$. Note that our indexes $a$ and $b$ start at 00 instead of $1$. In the rest of the section, we will adopt this convention.
If matrix $C_{\rho,l}$ is invertible, then we can have an unbiased estimation of $\Pr[z_{S}>0]$ using Algorithm 5, below. LABEL:prop:feasible, proved in Section 7.5, shows that $C_{\rho.l}$ is invertible for $\rho>1/2$ and thus completes the specification of our randomized response algorithm.
Algorithm 5 Unbiased estimate for number of reachable set $J_{m}(S)$
@@
Given any positive integer $l\geq 1$ and $\rho\in[0,1/2)$, matrix $C_{\rho,l}$ is invertible.
Note that when $\rho=1/2$, $\tilde{{\mathbf{x}}}$ consists of uniform random bits, and we cannot recover $\Pr[z_{S}>0]$ from the perturbed influence samples, and LABEL:prop:feasible shows that we can recover $\Pr[z_{S}>0]$ for any $\rho<1/2$.
@@
The above theorem, proved in Section 7.6, provides a tail bound on the approximation error of $\mathcal{M}_{loc}^{k,\epsilon}$ given a fixed number of influence samples. We can convert the above result to a high probability error bound by choosing $m=\Omega({k^{3}V_{\rho_{\epsilon},k}^{2}\ln n}/{\alpha^{2}})$ which for $k$ constant and $\epsilon\gg\alpha$ improves on the $1/\alpha^{3}$ sample complexity in LABEL:thm:original (complete proof in Section 7.7):
@@
Note that $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})$ in LABEL:thm:rr and LABEL:thm:rr:acc can be simplified as follows: $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})=O((1/2-1/(e^{\epsilon}+1))^{-k^{2}})=O([2(e^{\epsilon}+1)/(e^{\epsilon}-1)]^{k^{2}})$. Therefore $V_{\rho_{\epsilon},k}=O(\epsilon^{-k^{2}})$ as $\epsilon\to 0$ and $V_{\rho_{\epsilon},k}=O(2^{k^{2}})$ as $\epsilon\to\infty$.
@@ 4 The Nature of Our Privacy Protections in Social Networks
From the perspective of privacy guarantees, exponential mechanism (central DP) only protects the input data against the algorithmic revelations of the output (seeded nodes), giving the algorithm itself unhindered access to user data. On the other hand, the randomized response mechanism (local DP) provides a stronger privacy notion that protects the input data directly against any (adversarial) misuse, regardless of – and including — their use for seeding. Our results address both central and local notions of privacy, providing performance (utility) and sample complexity guarantees for influence maximization algorithms whose outputs/inputs are randomized at a desired level to give central/local DP protection given our input data structure, i.e., influence samples.
@@
In our framework, we explicitly establish a budget $(\epsilon)$ for privacy leaks through the observed network cascades. This budget is a quantifiable limit on how much privacy/security risk one is willing to tolerate across all queries, not just a single query of the data. The latter highlights yet another way, in which our formulation of privacy budget is distinct from the setting of node- and edge-DP for graph input data structures. While algorithmically tractable, practice-wise our notion of adjacency protects each data donor against revelations of their contributions to any given cascade. While a donor may be willing to donate their contagion history, there may be some cascades in that history that they feel uncomfortable about (revealing sensitive or compromising information). Or the donors may dynamically contribute their contagion data where previous donations are irrevocable. Our notion of influence sample differential privacy provides donors with plausible deniability at any fixed cascade. This way donors are incentivized to contribute their entire set of $m$ cascades because we can bound their privacy risks from information leakage at any specific cascade in the set of size $m$, either centrally or locally - depending on the data curation environment - using our Algorithms 3 and 4. As long as the number of sensitive cascades is much smaller than $m$, we believe this notion of adjacency provide meaningful privacy protections. This is reflected in the requisite randomization in our greedy exponential mechanism (Algorithm 3): to achieve $\epsilon-$DP, each node $v$ is chosen with probability proportional to $\exp\left(\frac{\epsilon m}{2n}I_{\mathbf{x}}(v|S)\right)$, where $I_{\mathbf{x}}(v|S)$ is the marginal benefit of adding $v$ to the current seed set $S$ and it is estimated from the input influence samples ${\mathbf{x}}$. As $m\to 0$, this “soft max” converges to a uniform random choice, completely ignoring the input influence samples ${\mathbf{x}}$. On the other hand, the randomized response mechanism achieves $\epsilon-$DP locally by flipping the status of nodes in each cascade independently at random with probability $1/(1+e^{\epsilon})$. While the randomized response noise level is determined independently of the sample size $m$, the latter scales up as $m=O({k^{3}\epsilon^{-2k^{2}}\ln n}/{\alpha^{2}})$ with the decreasing privacy budget $(\epsilon)$ to sustain algorithmic performance in the privacy-critical regime ($\epsilon\to 0$).
@@
As a concrete case, consider the choice of community leaders to undergo training for a public health or education intervention. To identify and target community leaders for treatment, the interventionist can survey the social network connections among the population. For example, Banerjee et al. (2013) survey individuals to collect social network data along 12 dimensions, including whom they get medical advice from or give medical advice to, relative and non-relatives with whom they socialize, and people with whom they exchange goods. While the community leaders may consent to and be properly compensated for their time in treatment, their social contacts face a privacy risk in having one of their associates chosen. As such, people may refrain from contributing to social network data collection or refuse to answer survey questions, especially when association with the treated individuals can pose harm, such as stigma or retaliation. For example, to prevent school bullying a few students may be recruited into treatment programs — e.g., by encouraging them to take a public stance against conflict (Paluck et al. 2016) — if the treated individuals are chosen based on their social network positions, and even if they are protected in treatment, their friends and contacts can face retaliation outside the treatment. Our central privacy protections in Section 3.1 allows the network participants to make a credible claim that their network associates are recruited into treatment by random choice, irrespective of their associations or contribution to the collected network cascade data.
@@
In other related contexts, the nature of collected data can itself be compromising, irrespectively of how individuals are selected into treatments; e.g., surveillance and control of sexually transmitted disease and HIV may entail surveys of sexual contacts or needle sharing networks among people who inject drugs (Fairley et al. 2022, Endo et al. 2022). Administering such privacy-critical data should entail local privacy protections, whereby individual data is randomized as they are collected before any intervention designs or allocation decisions. Our local privacy guarantees in Section 3.2 are suited for such cases where individual contagion data may be collected for intervention design, but needs to be protected locally at the collection phase.
@@
In a similar vein, social media platforms rely on social network data and information from past diffusions to organize their ad campaign and improve the efficacy of targeted advertising. Ensuring that these network operations do not violate users’ privacy is critical to the sustainability of social media platforms and their ad economies. Minority groups, in particular, can face more privacy risks than others in a social network. For example, a few adversarial nodes can coordinate their actions to move a web crawler towards or away from certain parts of an online social network to compromise their privacy through a re-identification attack (Backstrom et al. 2007). More broadly, deployment of learning systems to improve user experience requires accessing and aggregating historical usage information (e.g., for emojis and web domains) to optimize resource overhead for computation and communication, while respecting the personal and sensitive nature of the collected data (Apple Differential Privacy Team 2017, Erlingsson 2014b).
5 Simulation Results

### v3

@@ 3 Differentially Private Seeding Algorithms
In this section, we provide two DP seeding algorithms that satisfy the DP requirements either centrally or locally. First in Section 3.1, we use an exponential mechanism that ensures that the seed selection output is centrally DP. Next in Section 3.2, we use the randomized response algorithm to achieve local DP according to the criteria set forth in Definition 2.4. LABEL:thm:exp:acc and LABEL:thm:rr:acc provide our sample complexity bounds in each case to achieve $\epsilon$-ISDP with a $(1-1/e)\mbox{\sf OPT}-\alpha n$ performance guarantee. In particular, we observe that the sample complexity bound for achieving local DP is much harsher at $O(k^{3}\epsilon^{-2k^{2}}\ln n/\alpha^{2})$ compared to $O(k\ln n/(\alpha\epsilon))$ for central DP. This exponential separation of the dependency of sample complexity on $k$ is reflected in our simulation results in Section 5, where the exponential mechanism shows a faster improvement in performance with an increasing number of influence samples and its output has a significantly higher expected spread size when the privacy budget ($\epsilon$) is low.
@@ 3.1 Central DP with Exponential Mechanism
In this section, we design a DP seeding algorithm using the exponential mechanism and modifying Algorithm 1. To protect privacy, instead of Algorithm 2, we use an exponential mechanism that differentially outputs a softmax (LABEL:thm:exp:utility). Formally, given $\epsilon>0$ and the integer $k$ we define a seeding algorithm $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ that runs Algorithm 1 but replaces the single seed selection steps of Algorithm 2 by its variant of the exponential mechanism Algorithm 3 $\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}},S)$. Note that Algorithm 3 is an exponential mechanism that uses $I_{\mathbf{x}}(\cdot|S)$ as the scoring function.
@@
Note that a naive algorithm that runs the exponential mechanism directly on all $k$ seeds at once is not efficient because the size of the output space is combinatorially large $\Omega(n^{k})$. To ensure that our algorithm is efficient, we sequentially run the exponential mechanism $k$ times to estimate and choose seeds similar to the greedy algorithm of Eckles et al. (2022), but reducing their time complexity to $O(nmk)$. LABEL:thm:exp:inf shows that our modified greedy algorithm ($\mathcal{M}_{exp}^{k,\epsilon}({\mathbf{x}})$) that applies the exponential mechanism ($\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}},S)$) for single seed selection $k$ times, is $\epsilon$-ISDP following Definition 2.4, and also provides a tail bound on the approximation error for maximization of $I_{\mathbf{x}}$ using $\mathcal{M}_{exp}^{k,\epsilon}({\mathbf{x}})$. Its proof in Section 7.3 follows from the facts that $I_{\mathbf{x}}$ is a submodular function, and the exponential mechanism in Algorithm 3 can be seen as a noisy maximum function that implements an approximate greedy step in LABEL:thm:exp:utility.
@@
Given graph ${\mathcal{G}}$ on $n\geq 2$ nodes and $k\leq n$, for any $0<\alpha\leq 1$ and $\epsilon>0$, $\mathcal{M}^{k,\epsilon}_{exp}(\cdot)$ that runs Algorithm 1 but replaces the single seed selection steps of Algorithm 2 by its exponential mechanism variant in Algorithm 3, $\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}},S)$, is $\epsilon$-ISDP and its output infects at least $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes with high probability, using at least $m=\max\{\frac{12}{\alpha\epsilon},\frac{9}{\alpha^{2}}\}k\ln n$ influence samples in $O(knm)$ run time.
Note that if we are not concerned with privacy and set $\epsilon\gg\alpha$, then the number of samples to achieve the additive error $\alpha n$ is of the order $\alpha^{-2}k\ln n$, which improves the $k^{2}(\ln n/\alpha)/\alpha^{3}$ bound of Eckles et al. (2022). Moreover, our result does not violate Theorem 3 of Eckles et al. (2022) which shows that $\Omega(k^{2})$ influence samples are required to achieve the $(1-1/e)\mbox{\sf OPT}-\alpha n$ performance guarantee for all $\alpha<(1-1/e)/k$, because our algorithm also needs $\Omega(k^{2})$ samples when $\alpha<1/k$.
@@ 3.2 Local DP with Randomized Response
We can achieve a stronger notion of privacy, local DP, by adding noise to the input data before running our (or any other) algorithms. Line 3 of Algorithm 4 adds the necessary amount of noise to achieve local $\epsilon$-ISDP in the input dataset ${\mathbf{x}}$ — flipping each entry of ${\mathbf{x}}$ independently with probability $\rho_{\epsilon}=1/(1+e^{\epsilon})$ to generate noisy input $\tilde{{\mathbf{x}}}$. However, the added input noise will bias our estimate of each candidate seed’s quality; in particular, $\operatorname*{\mathbb{E}}_{\tilde{{\mathbf{x}}}}[I_{\tilde{{\mathbf{x}}}}(S|T)]$ is not necessarily equal to $I_{\mathcal{G}}(S|T)$ for a pair of subsets $S,T\subset V$ and we no longer have the benefit of applying Proposition 2.1 directly. Rather in Algorithm 5 we propose a post-processing of the noisy input data $\tilde{{\mathbf{x}}}$ to derive an unbiased estimate of a seed set’s quality. The importance of proper post-processing steps in the analysis of DP data is highlighted in other applications, e.g., in the release of census DP data, where ignoring them can lead to biased decision-making and suboptimal results (Abowd et al. 2019, Ashmead et al. 2019, Zhu et al. 2022, 2021, 2023). The added variability of the estimators due to the DP noise can also degrade our estimation performance (even if not biased), which is a second concern in local DP algorithm design and opens a venue to explore alternative post-processing methods with potentially superior bias-variance tradeoff. In the remainder of this section, we first introduce Algorithms 4 and 5 and then offer LABEL:thm:rr:acc for their privacy and utility guarantees, addressing both concerns about bias and variance of decision making with local DP influence samples.
@@
Given a set $S$, an influence sample $x\in\{0,1\}^{n}$, and its perturbed version $\tilde{x}$, consider two sets of nodes $S\cap x$ and $S\cap\tilde{x}$. By Proposition 2.1, the influence function $I_{\mathcal{G}}(S)$ can be estimated from the reachable set, $n\Pr[S\cap x\neq\emptyset]$. However, to achieve local DP, we can only observe the perturbed, reachable set $S\cap\tilde{x}$. In other words, let $\tilde{z}_{S}:=|S\cap\tilde{x}|$ and ${z}_{S}:=|S\cap{x}|$ be two random variables. If we can estimate $\Pr[z_{S}=0]$ from our observed samples of $\tilde{z}_{S}$, then we can estimate the size of the reachable set from $I_{\mathcal{G}}(S)=n\Pr[z_{S}>0]$ — using Proposition 2.1 — and then run the standard greedy algorithm for seed selection based on estimated values.
Now, we discuss how to estimate $\Pr[z_{S}=0]$ from samples of $\tilde{z}_{S}$. The main observation is that the probability mass function of $\tilde{z}_{S}$ is a linear function of the probability mass function of $z_{S}$. Note that $S\cap x$ is the set of elements in $S$ that is also in the influence sample $x$. Hence, $S\cap\tilde{x}$ can be partitioned into (1) the subset $(S\cap\tilde{x})\cap x$ of elements of $x$ that remain in $S$ after perturbation and (2) the subset $(S\cap\tilde{x})\setminus x$ of the elements of $\tilde{x}$ that join $S$ after perturbation. The size of the first subset is a binomial random variable $Bin(|S\cap x|,1-\rho)$, and the size of the second subset is $Bin(|S\setminus x|,\rho)$. Therefore, given $|S|=l$ for all $a$ and $b$, the probability of $\tilde{z}_{S}=b$ conditional on $z_{S}=a$ is $\Pr[\tilde{z}_{S}=a|z_{S}=b]=\Pr[Bin(b,1-\rho)+Bin(l-b,\rho)=a],$ for all $a,b=0,\dots,l$ where $Bin(b,1-\rho)$ and $Bin(l-b,\rho)$ are independent. Furthermore, if we set the matrix $C=C_{\rho,l}\in{\mathbb{R}}^{(l+1)\times(l+1)}$ as the likelihood function where $C(a,b)=C_{\rho,l}(a,b)=\Pr[\tilde{z}_{S}=a|z_{S}=b]$, then we have the following:
@@
We can write each entry of the $C_{\rho,l}$ matrix explicitly as follows:
@@
To see why, consider the event that $\ell$ of the $z_{S}=b$ nodes in $S\cap x$ are flipped, leaving the $b-\ell$ nodes in $S\cap x\cap\tilde{x}$. To obtain $\tilde{z}_{S}=a$ nodes in $S\cap\tilde{x}$, the remaining $a-b+\ell$ nodes should come from perturbing the status of the $l-b$ nodes belonging to $S\setminus x$. Note that our indices ($a$ and $b$) start at 00 instead of $1$. In the remainder of the section, we will adopt this convention.
If the matrix $C_{\rho,l}$ is invertible, then we can have an unbiased estimate of $\Pr[z_{S}>0]$ using Algorithm 5, below. LABEL:prop:feasible, proved in Section 7.5, shows that $C_{\rho,l}$ is invertible for $\rho>1/2$ and thus completes the specification of our randomized response algorithm.
Algorithm 5 Unbiased estimate for the size of the reachable set $J_{m}(S)$
@@
Given any positive integer $l\geq 1$ and $\rho\in[0,1/2)$, the matrix $C_{\rho,l}$ is invertible.
Note that when $\rho=1/2$, $\tilde{{\mathbf{x}}}$ consists of uniform random bits and we cannot recover $\Pr[z_{S}>0]$ from perturbed influence samples. LABEL:prop:feasible shows that we can recover $\Pr[z_{S}>0]$ for any $\rho<1/2$.
@@
The above theorem, proved in Section 7.6, provides a tail bound on the approximation error of $\mathcal{M}_{loc}^{k,\epsilon}$ given a fixed number of influence samples. We can convert the above result to a high probability error bound by choosing $m=\Omega({k^{3}V_{\rho_{\epsilon},k}^{2}\ln n}/{\alpha^{2}})$ which for $k$ constant and $\epsilon\gg\alpha$ improves the sample complexity ($1/\alpha^{3}$) in LABEL:thm:original (complete proof in Section 7.7):
@@
We conclude this section by considering two scenarios: the non-private regime $\epsilon\to\infty$ and the full-privacy regime $\epsilon\to 0$. First, in the non-private regime ($\epsilon\to\infty$), $\rho_{\epsilon}\to 0$, which causes the matrix $C_{\rho,l}$ and its inverse to converge to the identity matrix, so we can take $V_{\rho_{\epsilon},k}$ to be a constant arbitrarily close to one. As a result, $m=O\left(\frac{k^{3}\ln n}{\alpha^{2}}\right)$ samples are sufficient, which is comparable to LABEL:thm:original with an additional factor of $k$, but improves the dependence on $\alpha$. We may reduce the dependence on $k$ from $k^{3}$ to a smaller order, using techniques in adaptive data analysis Bassily et al. (2016). Informally, our analysis requires that the query error in any possible seed set $S$ with $|S|\leq k$ is small enough (Lemma 7.5) so that the output of the greedy algorithm is good enough. However, we may not need such stringent conditions, because the greedy algorithm does not necessarily query all possible seed sets. Alternatively, adaptive data analysis provides a procedure to support adaptive queries from the greedy algorithm with a better error bound.  [^1] Interestingly, the procedure in Bassily et al. (2016) adds Gaussian noise to each query and achieves a small error guarantee using techniques from an exponential mechanism similar to  LABEL:thm:exp:acc. Furthermore, in the non-private regime ($\epsilon\to\infty$), our exponential mechanism and the local DP seeding algorithm are identical, allowing us to use the analysis in LABEL:thm:exp:inf and derive an $O\left(\frac{k\ln n}{\alpha^{2}}\right)$ upper bound for the required number of influence samples. On the other hand, if $\epsilon\to 0$, $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})$ in LABEL:thm:rr and LABEL:thm:rr:acc can be simplified as follows: $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})=O((1/2-1/(e^{\epsilon}+1))^{-k^{2}})=O([2(e^{\epsilon}+1)/(e^{\epsilon}-1)]^{k^{2}})=O(\epsilon^{-k^{2}})$.
@@ 4 The Nature of Our Privacy Protections in Social Networks
From the perspective of privacy guarantees, the exponential mechanism (central DP) only protects the input data against the algorithmic revelations of the output (seeded nodes), giving the algorithm itself unhindered access to user data. On the other hand, the randomized response mechanism (local DP) provides a stronger privacy notion that protects the input data directly against any (adversarial) misuse, regardless of – and including – their use for seeding. Our results address both central and local notions of privacy, providing performance (utility) and sample complexity guarantees for influence maximization algorithms whose outputs/inputs are randomized at a desired level to give central/local DP protection given the input data structure of influence samples.
@@
In our framework, we explicitly establish a privacy budget $(\epsilon)$ to limit privacy leaks through observed influence cascades, such as those collected during contact tracing for HIV prevention. This budget quantifies the amount of privacy or security risk that can be tolerated across all queries, rather than focusing on a single query of the data. This approach distinguishes our privacy budget formulation from node and edge differential privacy (DP) commonly applied to graph data structures. From a privacy standpoint, our notion of adjacency protects each data donor by ensuring that their contribution to any given cascade remains private. Although a data donor may be willing to share their history of sexual activity as part of an ongoing contact tracing effort, there may be certain cascades in that history that they prefer to keep private. In addition, donors can dynamically contribute their contagion data knowing that prior contributions cannot be revoked. Our definition of influence sample differential privacy offers donors plausible deniability for any individual cascade. This ensures that they are incentivized to provide their entire set of $m$ cascades, knowing that their privacy risk in each cascade is bounded, centrally or locally, depending on the data curation environment. These privacy guarantees are implemented in our Algorithms 3 and 4, ensuring that even in sensitive settings such as HIV prevention, participants’ data is handled securely while allowing effective interventions. As long as the number of sensitive cascades is much smaller than $m$, this notion of adjacency provides meaningful central privacy protections. This is reflected in the randomization required in our greedy exponential mechanism (Algorithm 3): To achieve $\epsilon-$DP, each node $v$ is chosen with a probability proportional to $\exp\left(\frac{\epsilon m}{2n}I_{\mathbf{x}}(v|S)\right)$, where $I_{\mathbf{x}}(v|S)$ is the marginal benefit of adding $v$ to the current seed set $S$ and is estimated from the input influence samples ${\mathbf{x}}$. As $m\to 0$, this “softmax” converges to a uniform random choice, completely ignoring the input influence samples ${\mathbf{x}}$. On the other hand, the randomized response mechanism achieves $\epsilon-$DP locally by flipping the status of nodes in each cascade independently at random with probability $1/(1+e^{\epsilon})$. Although the level of randomized response noise is determined independently of the sample size $m$, the latter scales as $m=O({k^{3}\epsilon^{-2k^{2}}\ln n}/{\alpha^{2}})$ with the decreasing privacy budget $(\epsilon)$ to maintain algorithmic performance in the privacy-critical regime ($\epsilon\to 0$).
@@
As a concrete case, consider the choice of community leaders to undergo training for a public health or education intervention. To identify and target community leaders for treatment, the interventionist can survey social network connections in the population. For example, Banerjee et al. (2013) survey individuals to collect data about their social networks in 12 dimensions, including whom they receive medical advice from or give medical advice to, relative and non-relatives with whom they socialize, and people with whom they exchange goods. Although community leaders may consent to treatment and be compensated for their time, their social contacts face a privacy risk in their association with treated individuals. As such, people may refrain from contributing to the collection of social network data or refuse to answer survey questions, especially when association with treated individuals can pose harm, such as stigma or retaliation. For example, to prevent school bullying, a few students can be recruited into treatment programs, e.g., by encouraging them to take a public stance against conflict (Paluck et al. 2016). If treated individuals are chosen based on their social network positions and even if they are protected in treatment, their friends and contacts can face retaliation outside the treatment. Our central privacy protections in Section 3.1 allow network participants to make a credible claim that their network associates are recruited into treatment by random choice, regardless of their associations or contribution to the collected network cascade data.
@@
In other related contexts, the nature of the collected data can itself be compromising, regardless of how individuals are selected into treatments; e.g., surveillance and control of sexually transmitted diseases and HIV can involve surveys of sexual contacts or needle sharing networks among people who inject drugs (Fairley et al. 2022, Endo et al. 2022). Administering such privacy-critical data should entail local privacy protections, by which individual data are randomized as they are collected before any intervention designs or allocation decisions. Our local privacy guarantees in Section 3.2 are suitable for such cases where individual contagion data may be collected for intervention design, but must be protected locally in the collection phase.
@@
In a similar vein, social media platforms rely on social network data and information from previous diffusions to organize their ad campaign and improve the efficacy of targeted advertising. Ensuring that these network operations do not violate users’ privacy is critical to the sustainability of social media platforms and their ad economies. Minority groups, in particular, can face more privacy risks than others in a social network. For example, a few adversarial nodes can coordinate their actions to move a web crawler toward or away from certain parts of an online social network to compromise their privacy through a re-identification attack (Backstrom et al. 2007). More broadly, the deployment of learning systems to improve user experience requires accessing and aggregating historical usage information (e.g., emojis and web domains) to optimize resource overhead for computation and communication, while respecting the personal and sensitive nature of the collected data (Apple Differential Privacy Team 2017, Erlingsson 2014b).
5 Simulations with Synthetic and Empirically-Grounded Network Data Sets
We start with a sanity check using synthetic random graph models in Section 5.1 and then test our proposed framework for data collection and intervention design using highly sensitive data on sexual activity in an HIV prevention context in Section 5.2.
5.1 Simulations on Synthetic Random Graph Models
To test the performance of our algorithms in a sanitized environment, we generate an input using the Erdős-Rényi random graph model on $200$ nodes with edge probability $0.15$ and run the randomized response algorithm and the exponential mechanism with different privacy budgets ($\epsilon=0.1,0.5$, or $1$) and different numbers of influence samples ($m$ ranging from zero to $500$). The independent cascade probability is set to $0.03$. Figure 1 shows that the increase rate with $m$ slows with decreasing privacy budget for both the exponential mechanism and the randomized response algorithm.